# Tracailer: An Efficient Trajectory Planner for Tractor-Trailer Vehicles in Unstructured Environments

### v1

@@ Abstract
The tractor-trailer vehicle (robot) consists of a drivable tractor and one or more non-drivable trailers connected via hitches. Compared to typical car-like robots, the addition of trailers provides greater transportation capability. However, this also complicates motion planning due to the robot’s complex kinematics, high-dimensional state space, and deformable structure. To efficiently plan safe, time-optimal trajectories that adhere to the kinematic constraints of the robot and address the challenges posed by its unique features, this paper introduces a lightweight, compact, and high-order smooth trajectory representation for tractor-trailer robots. Based on it, we design an efficiently solvable spatio-temporal trajectory optimization problem. To deal with deformable structures, which leads to difficulties in collision avoidance, we fully leverage the collision-free regions of the environment, directly applying deformations to trajectories in continuous space. This approach not requires constructing safe regions from the environment using convex approximations through collision-free seed points before each optimization, avoiding the loss of the solution space, thus reducing the dependency of the optimization on initial values. Moreover, a multi-terminal fast path search algorithm is proposed to generate the initial values for optimization. Extensive simulation experiments demonstrate that our approach achieves several-fold improvements in efficiency compared to existing algorithms, while also ensuring lower curvature and trajectory duration. Real-world experiments involving the transportation, loading and unloading of goods in both indoor and outdoor scenarios further validate the effectiveness of our method. The source code is accessible at https://github.com/ZJU-FAST-Lab/tracailer/.
@@ I Introduction
In recent years, autonomous driving has gained a lot of interest and great growth due to its potential social benefits. While when large cargoes need to be transported on the ground, people will turn their attention to tractor-trailer systems, such as semi-trucks, as they can carry more via trailers. Tractor-trailer robots are a class of vehicles consisting of a driveable tractor and many unpowered trailers. The trailers are often connected to the tractor or other trailers by hinges, so that the motions of the trailers can be controlled by controlling that of the tractor[1].
Motion planning is a crucial component of autonomous navigation frameworks, and tractor-trailer robots are no exception. Similar to car-like robots, the trajectory planner for tractor-trailer robots aims to generate trajectories that minimize time consumption while satisfying kinematic constraints, enabling smooth and safe navigation to the target area. While motion planning for car-like robots is a relatively mature field, these techniques cannot be directly applied to tractor-trailer robots, since it is necessary to consider not only the state of the tractor but also the state of the trailer when moving. Specifically, compared to car-like robots, the difficulties in motion planning for tractor-trailer robots are mainly caused by the following three factors:
@@
    Complex kinematics: The hinge structure and the non-driveability of the trailers introduce more complex kinematic constraints. The coupling of the underactuated constraints of the trailers and the nonholonomic constraints of the tractor[2] requires the planner capable of handling more equation constraints. Meanwhile, it must avoid the jackknife, which will make the entire robot system uncontrollable[3].
@@
    High-dimensional state space: Each trailer attached to the robot introduces three additional dimensions to the robot in space, i.e., position and yaw angle, which can significantly increase the time consumed by search- or sampling-based path planning algorithms[4, 5]. Similarly, for optimization-based methods[2, 6, 7, 8, 9], the increase in dimension of the state leads to an increase in the dimension of the optimization variables, which may result in more memory consumption and slower convergence.
@@
To cope with the first two problems, we draw inspiration from the efficiency of polynomial trajectories optimization and differential flatness in motion planning for car-like robots[10, 11], proposing a new trajectory representation for tractor-trailer robots. We model the spatio-temporal trajectory optimization problem of tractor-trailer robots in continuous space, thus eliminating the discretization of the motion process and substantially improving the efficiency of the optimization. However, the singularity of differential flatness brought about by kinematics of the tractor may lead to a decrease in the optimization success rate and smoothness of the trajectory, which may make trajectory tracking difficult. Thus, in this work, we introduce slackened arc length to eliminate this singularity, improving the success rate of optimization. In addition, this approach makes the feasibility constraints easier to satisfy, leading to a further enhancement in efficiency.
@@
    Moreover, we build an autonomous navigation system for tractor-trailer robots, including a real-time replanning and mapping module, demonstrating the effectiveness and efficiency of the proposed method in extensive simulations and real-world experiments. The software is open-sourced to promote further research in this field.
@@ II Related Works
Motion planning for tractor-trailer robots can be mainly classified into geometry-based methods and optimization-based methods. Geometry-based methods only plan the geometric path for the robot and design the path tracking controller to track it. The geometric path is usually generated by search- or sampling-based methods. Search-based methods typically pre-construct a grid map or graph structure that represents the environment, and then perform graph searching within it to find paths [12]. For example, Liu et al. [13] utilized the conception of equivalent size for obstacles growing and robot shrinking, planning the path by applying a heuristic genetic algorithm. Sun et al. [14] built a global compound roadmap by combining the local regular roadmap with the universal probabilistic roadmap, lowering the complexity of the planning computation. To improve the efficiency and path quality, Ljungqvist et al. [4] generated a finite set of kinematically feasible motion primitives offline, searching solutions for a general 2-trailer system in a regular state lattice created by these motion primitives. They also proposed a novel parametrization of the reachable state space to make the graph-search problem tractable for real-time applications. Furthermore, by simplifying the design of motion primitives and utilizing reinforcement learning to obtain a well-defined heuristic function, Leu et al. presented an improved A-search guided tree[15] that allows quick off-lattice exploration to find a solution.
Sampling-based methods usually define a space for a path planning task. They maintain a tree structure and iteratively sample in this space, expanding tree nodes according to some criteria, ultimately extracting the trajectories from the tree containing the start and goal. By customizing the sampling space based on motion primitives instead of the space containing all $SE(2)$ states of vehicles in the tractor trailer robot, Cheng et al. [16], Lattarulo et al. [5] used the Rapidly-exploring Random Tree[17, 18] (RRT) algorithm to efficiently plan a feasible path that satisfied with the nonholonomic and mechanical constraints. Moreover, Manav et al. [19] proposed iterative analytical method, combining which with Closed-Loop Rapidly Exploring Random Tree (CL-RRT) approach in a cascade path planning, enabling the generation of both kinematically feasible and deterministic parking maneuvers with obstacle avoidance.
@@
Although geometry-based methods can work in various scenarios, these approach always search in a discrete space, where the optimality of the obtained solution is positively correlated with the search duration and the granularity of the space. When information of higher dimensions in the state of the robot is considered to find a more optimal solution, these methods need to explore larger space, whereupon the problem of combinatorial explosion is magnified, severely affecting the efficiency. To find better solutions in such a large space, optimization-based approaches emerged. These methods often model motion planning as an optimization problem and use numerical solvers to solve it, whose initial values often derived from geometric paths obtained by search- or sampling-based methods. Trajectory optimization typically explores optimal solutions in a continuous space, which can take into account information about higher-order states with guarantees of local optimality. Muralidhara et al. [7] modeled the motion planning problem of tractor-trailer robots as Optimal Control Problem (OCP), describing the trajectory by states in discrete timestamps, solving it by direct multiple shooting method. But in obstacle avoidance, this approach only considers not leaving the tracking line of the traffic lane, which cannot be directly applied to unstructured complex environments. Mohamed et al. [26] proposed an approach combining artificial potential fields and optimal control theory to achieve a more generalized obstacle avoidance for tractor-trailer robot. However, this method can easily fall into a local optimum when the environment is complex, resulting in the robot cannot reach the target region. Also using the potential field approach to represent the environment, Wang et al. [27] adopt path integral policy improvement to optimize the paths from geometry based methods, solving the problem of poor performance of graph search methods in narrow scenes the limitations by resolution and search space. Nevertheless, this approach is time-consuming, making it difficult to apply to tasks that require online replanning, such as exploring unknown environments and navigating through dynamically changing scenarios.
@@
The algorithm proposed in this paper is also an optimization-based methods. To plan efficiently, we do not adopt discrete states to represent trajectory and model the problem as an OCP, but propose a lightweight, compact and high-order smooth trajectory representation, reducing the dimension of optimization variables and simplifying the problem.
@@
In works[28, 29, 8], both obstacles and robots are represented by many circles , as this simple form can be easily modeled into OCP. However, real-world environments are complex that cannot be accurately modeled using only circles. To represent the environment more concisely and accurately, some works [30, 2, 31, 9, 32] use convex polygons to approximate obstacles. They also modeled the motion planning problem for tractor-trailer robots as OCP, considering safety constraints and duration in the optimization. Their methods require applying constraints to each obstacle and each vehicle of the robot at each moment. This feature tightly couples their computational complexity to the environment, limiting the efficiency of planning in obstacle-dense environments. To address this issue, some works[6, 33] shift to Safe Travel Corridors (STC) as obstacle avoidance constraints. The constraints imposed by this method are only related to the trajectory length and not to the density of obstacles. However, the classical safe corridors must be computed before trajectory optimization, since they are constructed based on the initial values of the optimization. For motion planning of a tractor-trailer robot, this will take a lot of time, especially if the number of trailers is large, since the safety corridors need to be built for each vehicle of the robot separately. The second disadvantage of the safe corridor is its high dependence on initial value, while a good one requires much more time due to the high state dimension and complex kinematics of the robot. Usually, the process of safe corridor generation selects some state points of a trajectory as seeds, which will affect the location and size of the generated safe corridors. Thus in some scenarios where it is more difficult to obtain a good initial value of the trajectory, e.g., the tractor-trailer robot needs to make a turn in a narrow environment, the feasible region constrained by the safe corridors may not be able to contain sufficiently good solutions. To reduce the dependency of the optimization on initial values, Li et al.[34] proposed a multi-stage method that only uses paths from A* as the initial value for trajectory optimization of tractor-trailer robots traversing a curvy tunnel, improving some of the efficiency for this task.
@@ III Planning Framework
Our goal is to plan a kinematically feasible and collision-free trajectory with the initial state of the robot $[x_{0}^{\text{init}},y_{0}^{\text{init}},v_{0}^{\text{init}},\theta_{0}^{\text{init}},\boldsymbol{\theta}^{\text{init}}]^{\text{T}}$ where $\boldsymbol{\theta}^{\text{init}}=[\theta_{1}^{\text{init}},\theta_{2}^{\text{init}},...,\theta_{\text{N}}^{\text{init}}]$ and information about the obstacles $\textbf{Info}_{c}$, such that the robot can follow this trajectory and eventually arrive within the assigned target region. In this work, point cloud are used to characterize the obstacles information as it can be directly acquired from LiDAR. Besides, we utilize a 2D convex polygon $\textbf{Info}_{e}$ to describe the target region, which is a more general condition compared to the work[6] and can also brings convenience for trajectory optimization.
@@ III-A Trajectory Parameterization
Differential flatness and polynomial representations have brought a significant improvement in the efficiency of trajectory generation for car-like robots[10]. The trajectory representation in this work is also inspired by these. For the robot like the one in Fig. 1, there is work[36] has demonstrated it is a differential flat system with $\boldsymbol{p}_{\text{N}}$ as flat output. However, the order of the derivatives involved in the derivation process of physical variables such as the longitudinal acceleration and steering angle of the tractor is positively correlated with the number of trailers. Thus the nonlinearity of the system will sharply strengthen as the number of trailers increases. Meanwhile, this method cannot be extended to more general tractor-trailer robot system, where the trailer $\boldsymbol{p}_{i+1}$ is not hooked directly to the center $\boldsymbol{p}_{i}$ of the previous one ($i=1,2,...,\text{N}$), but at a distance from this point.[1]
@@
where $\kappa(t)$ is the curvature of the trajectory. However, when the robot is stationary, i.e., there is a moment $t=t_{s}$, such that $v_{0}(t_{s})=0$, the above part of the equations will have a zero denominator, which is known as the singularity of differential flatness[11]. To avoid this problem, existing works usually add small positive constants to the denominator[10] or set a minimum velocity constraint[37]. These can lead to larger trajectory tracking errors, as well as making it difficult for the optimizer to converge to a solution that satisfies the kinematic constraints, especially the curvature constraint.
To solve this problem, inspired by the work[38] that utilizes arc length $s(t)$ to represent trajectories, we introduce slackened arc length trajectory $\mathfrak{s}(t)$ and a constraint to avoid singularities. According to the work[38], the arc length $s(t)$ is satisfied: $s(0)=0,\dot{s}(t)=\sqrt{\dot{x}_{0}^{2}(t)+\dot{y}_{0}^{2}(t)}=v_{0}(t)\geq 0$. Also, Eq.(1) and Eq.(2) are rewritten as:
@@
To slacken the constraint (20) and facilitate optimization, we use a new variable $\mathfrak{s}(t)$ to replace the arc length $s(t)$. In this paper, we define slackened arc length $\mathfrak{s}(t)$ to satisfy the constraints $\mathfrak{s}(t)\geq 0$ and $\dot{\mathfrak{s}}(t)\geq 0$. Meanwhile, we rewrite Eq.(1) and Eq.(2) as:
$\dot{x}_{0}(\mathfrak{s})\dot{\mathfrak{s}}(t)$ $=v_{0}(t)\cos\theta_{0}(t),$ (22)
$\dot{y}_{0}(\mathfrak{s})\dot{\mathfrak{s}}(t)$ $=v_{0}(t)\sin\theta_{0}(t).$ (23)
@@
Same as in the work[39], quintic piecewise polynomials with four times continuously differentiable at the segmented points are used to represent the trajectories of the robot. Introducing $\mathfrak{s}(t)$, each piece of trajectory is denoted as:
@@
where $j=1,2,...,\text{M};k=1,2,...,\Omega$ is the index of piecewise polynomial, $T_{*},*=\{p,\theta\}$ is the duration of a piece of the trajectory, $c_{*}\in\mathbb{R}^{6},*=\{x_{i},y_{j},\theta_{ki}\}$ is the coefficient of polynomial, $\gamma(*)=[1,*,*^{2},...,*^{5}]^{\text{T}},*=\{t,\mathfrak{s}\}$ is the natural base. Here, we set the trajectory of trailers have more pieces (i.e. $\Omega>\text{M}$) to better fit the kinematic constraints (6).

### v2

# Tracailer: An Efficient Trajectory Planner for Tractor-Trailer Robots in Unstructured Environments

@@ Abstract
The tractor-trailer robot consists of a drivable tractor and one or more non-drivable trailers connected via hitches. Compared to typical car-like robots, the addition of trailers provides greater transportation capability. However, this also complicates motion planning due to the robot’s complex kinematics, high-dimensional state space, and deformable structure. To efficiently plan safe, time-optimal trajectories that adhere to the kinematic constraints of the robot and address the challenges posed by its unique features, this paper introduces a lightweight, compact, and high-order smooth trajectory representation for tractor-trailer robots. Based on it, we design an efficiently solvable spatial-temporal trajectory optimization problem. To deal with deformable structures, which leads to difficulties in collision avoidance, we fully leverage the collision-free regions of the environment, directly applying deformations to trajectories in continuous space. This approach not requires constructing safe regions from the environment using convex approximations through collision-free seed points before each optimization, avoiding the loss of the solution space, thus reducing the dependency of the optimization on initial values. Moreover, a multi-terminal fast path search algorithm is proposed to generate the initial values for optimization. Extensive simulation experiments demonstrate that our approach achieves several-fold improvements in efficiency compared to existing algorithms, while also ensuring lower curvature and trajectory duration. Real-world experiments involving the transportation, loading and unloading of goods in both indoor and outdoor scenarios further validate the effectiveness of our method. The source code is accessible at https://github.com/Tracailer/Tracailer.
Index Terms:
Motion Planning, Trajectory Optimization, Tractor-trailer Robot, Nonholonomic Dynamics
Note to Practitioners—This paper addresses the challenges of motion planning for tractor-trailer robots, which are crucial in industries like logistics and agriculture. Our approach introduces a lightweight trajectory planning method that efficiently generates safe and time-optimal paths while considering the unique kinematic constraints of these vehicles. By utilizing the environment’s collision-free areas, we enhance safety and reduce reliance on initial conditions. While our method shows significant efficiency improvements in simulations and real-world tests, its effectiveness may vary in extremely dynamic environments. Practitioners can easily adopt our open-source code to integrate this planning technique into their operations, improving the efficiency and safety of tractor-trailer systems and paving the way for future advancements in autonomous transportation.
@@ I Introduction
In recent years, autonomous driving has gained a lot of interest and great growth due to its potential social benefits. When transporting large cargo, attention often shifts to tractor-trailer systems, such as semi-trucks, due to their capacity to carry more through the use of trailers. Tractor-trailer robot is a class of vehicles consisting of a drivable tractor and many unpowered trailers. The trailers are often connected to the tractor or other trailers by hinges, so that the motions of the trailers can be controlled by controlling that of the tractor[2].
Motion planning is a crucial component of autonomous navigation frameworks, and tractor-trailer robots are no exception. Similarly to car-like robots, the trajectory planner for tractor-trailer robots aims to generate trajectories that minimize time consumption while satisfying kinematic constraints, enabling smooth and safe navigation to the target area. Although motion planning for car-like robots has been studied relatively extensively, these techniques cannot be directly applied to tractor-trailer robots, since it is necessary to consider not only the state of the tractor but also the state of the trailer when moving. Specifically, compared to car-like robots, the challenges in motion planning for tractor-trailer robots are mainly caused by the following three factors:
@@
    Complex kinematics: The hinge structure and the non-drivability of the trailers introduce more complex kinematic constraints. The coupling of the underactuated constraints of the trailers and the nonholonomic constraints of the tractor[3] requires the planner capable of handling more equation constraints. Meanwhile, it must avoid the jackknife, which will make the entire robot system uncontrollable[4].
@@
    High-dimensional state space: Each trailer attached to the robot introduces three additional dimensions to the robot in space, i.e. position and yaw angle, which can significantly increase the time consumed by search- or sampling-based path planning algorithms[5, 6]. Similarly, for optimization-based methods[3, 7, 8, 9, 10], the increase in dimension of the state leads to an increase in the dimension of the optimization variables, which may result in more memory consumption and slower convergence.
@@
To cope with the first two problems, we draw inspiration from the efficiency of polynomial trajectories optimization and differential flatness[11] in motion planning for car-like robots[12, 13], proposing a new trajectory representation for tractor-trailer robots. We model the spatial-temporal trajectory optimization problem of tractor-trailer robots in continuous space, thus eliminating the discretization of the motion process and substantially improving the efficiency of the optimization. Benefiting from differential flatness[11], we can explicitly derive higher-order states of the robot and reduce the dimensionality of the optimization variables. However, the singularity of differential flatness brought about by kinematics of the tractor may lead to a decrease in the optimization success rate and smoothness of the trajectory, which may make trajectory tracking difficult. Thus, in this work, we introduce slackened arc length to eliminate this singularity, improving the success rate of optimization. In addition, this approach makes the feasibility constraints easier to satisfy, leading to a further enhancement in efficiency.
@@
    Moreover, we build an autonomous navigation system for tractor-trailer robots, including a real-time replanning and mapping module, demonstrating the effectiveness and efficiency of the proposed method in extensive simulations and real-world experiments, such as those shown in Fig. LABEL:fig:head_image. The software is open-sourced to promote further research in this field.
@@ II Related Works
Motion planning for tractor-trailer robots can be mainly classified into geometry-based methods and optimization-based methods.
II-A Geometry-based Methods
Geometry-based methods usually plan only the geometric path and design the path tracking controller to track it. The geometric path is usually generated by search- or sampling-based methods. Search-based methods typically pre-construct a grid map or graph structure that represents the environment, and then perform graph searching within it to find paths [14]. For example, Liu et al. [15] utilized the conception of equivalent size for obstacles growing and robot shrinking, planning the path by applying a heuristic genetic algorithm. Sun et al. [16] built a global compound roadmap by combining the local regular roadmap with the universal probabilistic roadmap, lowering the complexity of the planning computation. To improve the efficiency and path quality, Ljungqvist et al. [5] generated a finite set of kinematically feasible motion primitives offline, searching solutions for a general 2-trailer system in a regular state lattice created by these motion primitives. They also proposed a novel parametrization of the reachable state space to make the graph-search problem tractable for real-time applications. Furthermore, by simplifying the design of motion primitives and utilizing reinforcement learning to obtain a well-defined heuristic function, Leu et al. presented an improved A-search guided tree[17] that allows quick off-lattice exploration to find a solution.
Sampling-based methods usually define a space for a path planning task. They maintain a tree structure and iteratively sample in this space, expanding tree nodes according to some criteria, ultimately extracting the trajectories from the tree containing the start and goal. By customizing the sampling space based on motion primitives instead of the space containing all $SE(2)$ states of vehicles in the tractor trailer robot, Cheng et al. [18], Lattarulo et al. [6] used the Rapidly-exploring Random Tree[19, 20] (RRT) algorithm to efficiently plan a feasible path that is satisfied with the nonholonomic and mechanical constraints. Moreover, Manav et al. [21] proposed an iterative analytical method, combining it with the Closed-Loop Rapidly Exploring Random Tree (CL-RRT) approach in cascade path planning, allowing the generation of both kinematically feasible and deterministic parking maneuvers with obstacle avoidance.
@@ II-B Optimization-based Methods
Although geometry-based methods can work in various scenarios, these approach always search in a discrete space, where the optimality of the solution obtained is positively correlated with the duration of the search and the granularity of the space. When information of higher dimensions in the state of the robot is considered to find a more optimal solution, these methods need to explore larger space, whereupon the problem of combinatorial explosion is magnified, severely affecting the efficiency. To find better solutions in such a large space, optimization-based approaches emerged. These methods often model motion planning as an optimization problem and use numerical solvers to solve it, whose initial values often derived from geometric paths obtained by search- or sampling-based methods. Trajectory optimization typically explores optimal solutions in a continuous space, which can take into account information about higher-order states with guarantees of local optimality. Muralidhara et al. [8] modeled the motion planning problem of tractor-trailer robots as Optimal Control Problem (OCP), describing the trajectory by states in discrete timestamps, solving it by direct multiple shooting method. But in obstacle avoidance, this approach only considers not leaving the tracking line of the traffic lane, which cannot be directly applied to unstructured complex environments. Mohamed et al. [28] proposed an approach combining artificial potential fields and optimal control theory to achieve a more generalized obstacle avoidance for tractor-trailer robot. However, this method can easily fall into a local optimum when the environment is complex, preventing the robot from reaching the target region. Also using the potential field approach to represent the environment, Wang et al. [29] adopt path integral policy improvement to optimize the paths from geometry based methods, solving the problem of poor performance of graph search methods in narrow scenes the limitations by resolution and search space. Nevertheless, this approach is time-consuming, making it difficult to apply to tasks that require online replanning, such as exploring unknown environments and navigating through dynamically changing scenarios.
@@
The method proposed in this paper is also an optimization-based algorithm. To plan efficiently, we do not adopt discrete states to represent trajectory and model the problem as an OCP, but propose a lightweight, compact and high-order smooth trajectory representation, reducing the dimension of optimization variables and simplifying the problem.
@@
In works[30, 31, 9], both obstacles and robots are represented by many circles, as this simple form can be easily modeled into OCP. However, real-world environments are complex and cannot be accurately modeled using only circles. To represent the environment more concisely and accurately, some works [32, 3, 33, 10, 34] use convex polygons to approximate obstacles. They also modeled the motion planning problem for tractor-trailer robots as OCP, considering safety constraints and duration in the optimization. Their methods require applying constraints to each obstacle and each vehicle of the robot at each moment. This feature tightly couples their computational complexity to the environment, limiting the efficiency of planning in obstacle-dense environments. To address this issue, some works[7, 35] shift to Safe Travel Corridors (STC) as obstacle avoidance constraints. The constraints imposed by this method are only related to the trajectory length and not to the density of obstacles. However, the classical safe corridors must be computed before trajectory optimization, since they are constructed based on the initial values of the optimization. For motion planning of a tractor-trailer robot, this will take a lot of time, especially if the number of trailers is large, since the safety corridors need to be built for each vehicle of the robot separately. The second disadvantage of the safe corridor is its high dependence on initial value, while a good one requires much more time due to the high state dimension and complex kinematics of the robot. Usually, the process of safe corridor generation selects some state points of a trajectory as seeds, which will affect the location and size of the generated safe corridors. Thus in some scenarios where it is more difficult to obtain a good initial value of the trajectory, e.g., the tractor-trailer robot needs to make a turn in a narrow environment, the feasible region constrained by the safe corridors may not be able to contain sufficiently good solutions. To reduce the dependency of the optimization on initial values, Li et al.[36] proposed a multi-stage method that only uses paths from A* as the initial value for trajectory optimization of tractor-trailer robots traversing a curvy tunnel, improving some of the efficiency for this task.
@@ III Planning Framework
Our goal is to plan a kinematically feasible and collision-free trajectory with the initial state of the robot $x_{0}^{\text{init}},y_{0}^{\text{init}},v_{0}^{\text{init}},\theta_{0}^{\text{init}},\boldsymbol{\theta}^{\text{init}}]^{\text{T}}$ where $\boldsymbol{\theta}^{\text{init}}=[\theta_{1}^{\text{init}},\theta_{2}^{\text{init}},...,\theta_{\text{N}}^{\text{init}}]$ and information about the obstacles $\textbf{Info}_{c}$, such that the robot can follow this trajectory and eventually arrive within the assigned target region. In this work, point cloud are used to characterize the obstacles information as it can be directly acquired from LiDAR. Besides, we utilize a 2D convex polygon $\textbf{Info}_{e}$ to describe the target region, which is a more general condition compared to the work[7] and can also brings convenience for trajectory optimization.
@@ III-A Trajectory Parameterization
Differential flatness and polynomial representations have brought a significant improvement in the efficiency of trajectory generation for car-like robots[12]. The trajectory representation in this work is also inspired by these. For the robot like the one in Fig. 1, Deligiannis et al.[38] have demonstrated it is a differential flat system with $\boldsymbol{p}_{\text{N}}$ as flat output. However, the order of the derivatives involved in the derivation process of physical variables such as the longitudinal acceleration and steering angle of the tractor is positively correlated with the number of trailers. Thus the nonlinearity of the system will sharply strengthen as the number of trailers increases. Meanwhile, this method cannot be extended to more general tractor-trailer robot system, where the trailer $\boldsymbol{p}_{i+1}$ is not hooked directly to the center $\boldsymbol{p}_{i}$ of the previous one ($i=1,2,...,\text{N}$), but at a distance from this point.[2]
@@
where $\kappa(t)$ is the curvature of the trajectory. However, when the robot is stationary, i.e., there is a moment $t=t_{s}$, such that $v_{0}(t_{s})=0$, the above part of the equations will have a zero denominator, which is known as the singularity of differential flatness[13]. To avoid this problem, existing works usually add small positive constants to the denominator[12] or set a minimum velocity constraint[39]. These can lead to larger trajectory tracking errors, as well as making it difficult for the optimizer to converge to a solution that satisfies the kinematic constraints, especially the curvature constraint. To demonstrate this point, we introduce a model predictive control (MPC)-based controller to measure the tracking error. Details can be found in Sec.IV-C and Sec.IV-D.
To solve this problem, inspired by the work[40] that utilizes arc length $s(t)$ to represent trajectories, we introduce slackened arc length trajectory $\mathfrak{s}(t)$ and a constraint to avoid singularities. Where $s$ and $\mathfrak{s}$ are both functions of the time $t$. For brevity, In the rest of the paper, we will simplify them to writing $s$ and $\mathfrak{s}$ instead of $s(t)$ and $\mathfrak{s}(t)$ when necessary. According to the work[40], the arc length $s(t)$ is satisfied: $s(0)=0,\dot{s}(t)=\sqrt{\dot{x}_{0}^{2}(t)+\dot{y}_{0}^{2}(t)}=v_{0}(t)\geq 0$. Also, Eq.(1) and Eq.(2) are rewritten as:
@@
To slacken the constraint (20) and facilitate optimization, we define a new variable slackened arc length $\mathfrak{s}(t)$ to replace the arc length $s(t)$. In this paper, we follow a part of the definition of the arc length $s(t)$ by setting $\mathfrak{s}(0)=0$ and $\dot{\mathfrak{s}}(t)\geq 0$. Similarly, we can rewrite Eq.(1) and Eq.(2) as:
$\dot{x}_{\text{new}}(\mathfrak{s}(t))\dot{\mathfrak{s}}(t)$ $=v_{0}(t)\cos\theta_{0}(t),$ (22)
$\dot{y}_{\text{new}}(\mathfrak{s}(t))\dot{\mathfrak{s}}(t)$ $=v_{0}(t)\sin\theta_{0}(t),$ (23)
Where both $x_{\text{new}}$ and $y_{\text{new}}$ are differentiable functions of $\mathfrak{s}$, $\mathfrak{s}$ is a differentiable function of time $t$, and $x_{\text{new}}(\mathfrak{s}(t))=x_{0}(t),y_{\text{new}}(\mathfrak{s}(t))=y_{0}(t)$. Note that Eq.(22) and Eq.(23) can be easily proved to be equivalent to Eq.(1) and Eq.(2), respectively, by the chain rule. Thus, for brevity, we will still use the symbols $x_{0}(\mathfrak{s}(t))$ and $y_{0}(\mathfrak{s}(t))$ instead of the symbols $x_{\text{new}}(\mathfrak{s}(t))$ and $y_{\text{new}}(\mathfrak{s}(t))$ in the subsequent parts of the paper.
@@
Same as in the works[41, 12] to consider human comfort by minimizing integration of jerk[42], based on the optimality conditions proved by Wang et al.[43] for the multi-stage control effort minimization problem, we choose quintic piecewise polynomials with four times continuously differentiable at the segmented points as the representation for trajectories of the robot. Introducing $\mathfrak{s}(t)$, each piece of trajectory is denoted as:
@@
where $j=1,2,...,\text{M};k=1,2,...,\Omega$ is the index of piecewise polynomial, $T_{*},*=\{p,\theta\}$ is the duration of a piece of the trajectory, $c_{*}\in\mathbb{R}^{6},*=\{x_{j},y_{j},\mathfrak{s}_{j},\theta_{ki}\}$ is the coefficient of polynomial, $\gamma(*)=[1,*,*^{2},...,*^{5}]^{\text{T}},*=\{t,\mathfrak{s}\}$ is the natural base.